OpenStudy (anonymous):

Define the region \(\Sigma\subset\mathbb{R^2}\) as a system of inequalities for \((x, y)\). \(\Sigma\) is the intersection between the two regions defined by the ellipses (1) and (2) - see below for equations. Assume that the ellipses intersect.

OpenStudy (anonymous):

\[\frac{y^2}{y_1^2}+\frac{x^2}{x_1^2}=1\]\[\frac{y^2}{y_2^2}+\frac{(x-x_0)^2}{x_2^2}=1\]As stated in assumptions, \(x_0>x_1-x_2\), and the resulting inequalities may be expressed in terms of all subscripted constants.

OpenStudy (anonymous):

It's also okay to express the inequalities for x in terms of y, seeing as this is a type II region.

OpenStudy (anonymous):

@lgbasallote @Hero @jim_thompson5910 @dumbcow Anybody mind helping please? No one's answered for 2 hours.

OpenStudy (anonymous):

@hero is here to save the day

hero (hero):

count me out

OpenStudy (experimentx):

x_0 is a point, x_1 - x_2 is a distance

OpenStudy (anonymous):

@Geometry_Hater

OpenStudy (anonymous):

@experimentX x_0, x_1, x_2, y_1, y_2 are all scalars

OpenStudy (anonymous):

okay, I guess I'm going to Math.SE again!

OpenStudy (dumbcow):

i would take each ellipse equation and solve for "y^2" then set them equal in order to get it in terms of x rearrange terms and finally solve for x_0 ...it gets messy with so many constants then substitute that expression into inequality .... x_0 > x_1 -x_2 then rearrange terms again to solve for "x" using quadratic formula

OpenStudy (anonymous):

Yes that would get us the two points of intersection, I agree. I can do that. But that still doesn't define the area.

OpenStudy (dumbcow):

what do you mean? you have to find the area of the region

OpenStudy (dumbcow):

i thought you had to define the region with inequalities... a<x<b and c<y<d is that right

OpenStudy (anonymous):

Yes. But say I have to elipses:  I admit that we can find a, b, and c. But how does that help us find the region in the middle?